A hydrocarbon,W with a mass 10.50 g has 17.35% hydrogen. W

undergoes complete combustion.


Calculate the mass (g) of H 2 O formed.


1
Expert's answer
2021-09-04T02:28:44-0400

Solution:

hydrocarbon = CxHy

Mass of CxHy = Mass of C + Mass of H


Mass of H = %H × Mass of CxHy

Mass of H = 0.1735 × 10.50 g = 1.82175 g H


The molar mass of H is 1.00784 g/mol.

Hence,

(1.82175 g H) × (1 mol H / 1.00784 g H) = 1.8076 mol H


2yH + yO → yH2O

Hence,

(1.8076 mol H) × (1 mol H2O / 2 mol H) = 0.9038 mol H2O


The molar mass of H2O is 18.01528 g/mol.

Hence,

(0.9038 mol H2O) × (18.01528 g H2O / 1 mol H2O) = 16.316 g H2O = 16.32 g H2O

Mass of H2O = 16.32 g


Answer: 16.32 grams of H2O formed.